the perimeter of a California King sized bed is 312 inches. the length is a foot longer than the width. find the dimensions of the bed.
Answer by stanbon(75887)

the perimeter of a California King sized bed is 312 inches. the length is a foot longer than the width. find the dimensions of the bed.
-----------
P = 2(L + W)
----
312 = 2((x+1)+x)
156 = 2x+1
2x = 155
x = 77.5 inches (width)
x+1 = 78.5 inches (length)
============
Cheers,
stan H.
============